OpenStudy (anonymous):

Joy added two and one-half cups of broth to her pot of soup. If the result was six and seven-eighths cups of soup, how much was in the pot to begin with? four and three-eighths cups nine and three-eighths cups four and five-eighths cups nine and five-eighths cups

OpenStudy (anonymous):

@Chineseboy15

OpenStudy (anonymous):

x + 2 1/2 = 6 7/8 (turn to improper fractions) x + 5/2 = 55/8 (multiply by the LCD, which is 8) 8x + 20 = 55 (subtract 20 from both sides) 8x = 55 - 20 8x = 35 x = 35/8 = 4 3/8

OpenStudy (anonymous):

any questions ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

good to hear :)